Question 8 Find the Rel(-1+-7i)/(-4+5i)) This is division. Question 9 Find the Im((9+-61)(-8+7i))
Answer #1

0 let z = 2 = (-2+72) (-4+ si) (-2+71) (-4-si ) (-4+5) (-4-si) (+71) (-4-5i) (-+)?-151)? 6 2 = (a-6) (a+b) (1. os restos z z2 let z = (9+61) (-a+71) 2= -72+631-481+4212 z = -727151 +42(-1) [::.=-1] =-724151-42 2 z= atib z =-114 tisi Re(2) Im (2) Im